solve the system of linear equations.

x-y + 2z = 5

2x - 2y + 4z = 10

3x - 3y + 6z = 15


1
Expert's answer
2021-03-30T16:28:18-0400

"x-y+2z=5\\
2x-2y+4z=10\\
3x-3y+6z=15""\Darr""x-y+2z=5\\
2x-2y+4z=10\space|\space/2\\
3x-3y+6z=15\space|\space/3""\Darr""x-y+2z=5\\
x-y+2z=5\\
x-y+2z=5""\Darr""x-y+2z=5"


Answer: ("y-2z+5"; "y"; "z") where "y\space\epsilon\spaceℝ", "z\space\epsilon\spaceℝ".